OpenStudy (anonymous):

The area of a rectangle is the product of width and length. If a rectangle has length 12mthe square root of 21 times m. and width the square root of 16 times m., which of the following is the area of the rectangle? 192mthe square root of 21 times m. 192m2the square root of 21. 24mthe square root of 21 times m. 48m2the square root of 21.

OpenStudy (anonymous):

length \(12m\sqrt{21m}\) and width \(\sqrt{16m}\) ? ?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

well, what is \(\sqrt{16}\) ?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

the answer is d